two towns A and B are 80km apart. A passenger bus travels daily from A to B in heavy mist it has to travel 20km/h slower than in good weather and reaches its destination 12 minutes later than the usual what is the speed of the bus in good weather.
Answer by lwsshak3(11628)

two towns A and B are 80km apart. A passenger bus travels daily from A to B in heavy mist it has to travel 20km/h slower than in good weather and reaches its destination 12 minutes later than the usual what is the speed of the bus in good weather.
***
let x=speed of bus in good weather
x-20=speed of bus in heavy mist
distance/speed=travel time
12 min=1/5 hr
..
80/(x-20)-80/x=1/5
lcd:x(x-20)
80x-80x+1600=(x^2-20x)/5
x^2-20x-8000=0
(x+80)(x-100)=0
x=100
speed of bus in good weather=100 km/hr